OpenStudy (anonymous):

Find the first, fourth, and tenth terms of the arithmetic sequence described by the given rule. a(n) -6 + (n - 1) (6)

OpenStudy (anonymous):

0, -18 and -54

OpenStudy (anonymous):

dude there is no -18 -54 as an answer only thing near that is 18 54

OpenStudy (welshfella):

please explain

OpenStudy (anonymous):

–6, 12, 48 –6, 18, 54 0, 18, 54 6, 18, 54 thats the possable outcomes

OpenStudy (welshfella):

plug in n = 1 and see what you get

OpenStudy (anonymous):

then you do 1-1 so thats 0

OpenStudy (welshfella):

a1 = -6 + (1-1)d

OpenStudy (welshfella):

* thats 6 not d

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

ok so the correct answer is D?

OpenStudy (welshfella):

so whata the first term?

OpenStudy (anonymous):

dude i really think my answer is correct you just do the previous answer -6 every time n goes up 1

OpenStudy (welshfella):

first term = -6 + 0*6 = -6

OpenStudy (anonymous):

Wait the correct one is B

OpenStudy (anonymous):

and what was answer B?

OpenStudy (welshfella):

second = -6 + (2-1)*6 = ?

OpenStudy (welshfella):

4th term = -6 +3*6 = 12

OpenStudy (anonymous):

ok

OpenStudy (welshfella):

10th = -6 + (10-1)*6 = -6 + 9*6 = ?

OpenStudy (anonymous):

54

OpenStudy (welshfella):

No -6 + 54 = ?

OpenStudy (anonymous):

48

OpenStudy (welshfella):

right A is the correct option

OpenStudy (anonymous):

thanks so much i fanned you:)

OpenStudy (welshfella):

ty yw